OpenStudy (anonymous):

A prism has total surface area of 360 m2 and volume of 60 m3. If the length, width, and height are reduced to half their original sizes, what will be the following? a. the new surface area b. the new volume Solve for the following if the height is tripled. c. the new surface area d. the new volume

OpenStudy (anonymous):

\[S' = \frac{ Width }{ 2 } \times \frac{ Length }{ 2 } \times \frac{ Height }{ 2 } = \frac{ S }{ 8 }\] \[A' = \frac{ Width }{ 2 } \times \frac{ Length }{ 2 } = \frac{ A }{ ? }\]

whitemonsterbunny17 (whitemonsterbunny17):

Oh gosh, Math... Not my forte... Maybe @Astrophysics or @sleepyjess can help. :)

OpenStudy (anonymous):

Thanks

OpenStudy (sleepyjess):

Dumb question here, but my mind is blanking, what is S' and A'?

OpenStudy (anonymous):

Surface and Area

OpenStudy (sleepyjess):

Surface area of a prism is \(SA = 2*BA + LA\)

OpenStudy (sleepyjess):

I'm drawing a blank... maybe @zepdrix ?